what is the answer to this problem?
2x+5y=-16
y=-20-6x

Answer by fractalier(6550)

Let us use the substitution method for this one...from
2x+5y=-16
y=-20-6x
we can substitute -20-6x in for "y" in the first equation and get
2x + 5(-20-6x) = -16
2x - 100 - 30x = -16
-28x - 100 = -16
-28x = 84
x = -3
If x = -3, we can substitute into the second equation and get y...
y = -20 - 6(-3)
y = -20 + 18
y = -2